OpenStudy (anonymous):

Solve. 2x2 + 12x= 10

OpenStudy (mertsj):

Divide both sides by 2: x^2+6x = 5 Complete the square by adding 9 to both sides. x^2+6x+9 = 14 Factor the left side. (x+3)^2=14 Take the square root of both sides. x+3= +/- sqrt14 Subtract 3 from both sides: x=-3+/-sqrt14

OpenStudy (anonymous):

Thank you so much!! :)